# Connectivity of partial basis complexes of freely decomposable groups

### v1

@@ 2.3. Posets of core graphs
This is because every $H\in\operatorname{X}(G,l)$ contains a unique maximal core subgraph $\mathring{H}\in\operatorname{C}(G,l)$. The assignment $H\mapsto\mathring{H}$ defines a monotone poset map $\operatorname{X}(G,l)\to\operatorname{C}(G,l)$ such that $H\geq\mathring{H}$ and restricts to the identity on $\operatorname{C}(G,l)$. Hence $\operatorname{X}(G,l)$ deformation retracts to $\operatorname{C}(G,l)$ (see [Bru:buildingsfreefactora, Corollary 2.5]). ∎
In [Bru:buildingsfreefactora, Proposition 4.14], it was shown that for $k(l)=0$, the poset $\operatorname{X}(G,l)$ is contractible if and only if $G$ has a separating edge. The following definition is set up in such a way that this result generalises to the cases $k(l)>0$.
@@
Note that for $k\leq 1$, an edge is $l$-separating if and only if it is separating and $G$ has non-trivial fundamental group. In particular, the $l$-separating edges of $(G,l)$ are the same as the $\emptyset$-separating edges of the trivially labelled graph $(G,\emptyset)$.
@@
To explain the case distinction, note that if $k(l)=1$, then $\operatorname{X}(G,l)=\operatorname{X}(G,\emptyset)$. In [Bru:buildingsfreefactora, Proposition 4.14], Lemma 2.3 was deduced from [Vog:Localstructuresome, Proposition 2.2] for the case $k=0$, $n\geq 2$ under the assumption that every vertex of $G$ has valence at least three.
Proof of Lemma 2.3
Suppose first that $G$ contains an $l$-separating edge $e$. Then by Item 3 of Definition 2.2, $G\setminus e$ is an element of $\operatorname{X}(G,l)$. Item 2 of Definition 2.2 implies that for all $H\in\operatorname{X}(G,l)$, also $H\setminus\left\{e\right\}$ is an element of $\operatorname{X}(G,l)$. The assignment $H\mapsto H\setminus\left\{e\right\}$ gives a monotone poset map $\operatorname{X}(G,l)\to\operatorname{X}(G,l)$, so a homotopy equivalence to its image. The image has a maximal element given by $G\setminus\left\{e\right\}$, so it is contractible. Hence, also $\operatorname{X}(G,l)$ is contractible.
Now assume that $G$ contains no $l$-separating edges. We prove that $\operatorname{X}(G,l)$ is a non-trivial wedge of $d(G,l)$-spheres by induction on the number of edges of $G$. If $G$ has just a single edge, then this edge is not $l$-separating and $\operatorname{X}(G,l)$ is empty, so in particular non-contractible. (This is true independently of whether the edge forms a loop and whether it has labelled endpoints or not.) Hence, the claim is true. In what follows, we will assume that there are at least two edges.
@@
Suppose that $G$ contains a loop $e$. Then the assumption that $G$ has at least two edges and that none of them is $l$-separating implies that $G\setminus e$ is an element of $\operatorname{X}(G,l)$. The poset $\operatorname{X}(G,l)$ can be written as the union of two subposets. Let
@@
where $l^{\prime}$ is the labelling induced by $l$. This subposet has $G\setminus e$ as its unique maximal element and hence is contractible. The second subposet is given by $\operatorname{X}(G\setminus e,l^{\prime})$ together with all proper subgraphs of $G$ containing $e$:
@@
By the previous paragraph, we now can assume that $G$ contains no loops. Take an edge $e$ which is either non-separating (and not a loop) or a leaf. Let $Y_{1}:=\operatorname{X}(G,l)\setminus\left\{G\setminus e\right\}$ and denote by $(G\setminus e,l^{\prime})$ the labelled graph obtained from $G$ by removing $e$.
@@
To prove this, we consider two cases. The first is that $e$ has both endpoints labelled. As $G$ has more than one edge, this implies that $\left\{e\right\}\in Y_{1}$. We get a monotone map $Y_{1}\to Y_{1}$ by the assignment $H\mapsto H\cup\left\{e\right\}$. Its image has a unique minimal element $\left\{e\right\}$, so $Y_{1}$ is contractible. Furthermore, as $G$ has no $l$-separating edge, the graph $G\setminus e$ is an element of $\operatorname{X}(G,l)$, i.e. $Y_{1}\neq\operatorname{X}(G,l)$, and $G\setminus e$ has no $l^{\prime}$-separating edge.
The second case is that at least one of the endpoints of $e$ is unlabelled. In this case, for all $H\in Y_{1}$, the image of $H$ in $G/e$ is also an element of $\operatorname{X}(G/e,\tilde{l})$, where $\tilde{l}$ is the labelling induced by $l$. This gives a natural poset map
$f\colon Y_{1}$ $\to\operatorname{X}(G/e,\tilde{l})$
that sends a subgraph to its image in the quotient $G/e$. We define a poset map in the other direction as follows: Let $v_{e}$ be the vertex of $G/e$ to which $e$ was collapsed. Then define
$g\colon\operatorname{X}(G/e,\tilde{l})$ $\to Y_{1}$
$K$$\mapsto\begin{cases}K\cup\left\{e\right\}&v_{e}\in K,\\ K&\text{otherwise}.\end{cases}$
Here we have identified $K$ with the set of corresponding edges in $G$. We have $g\circ f(H)\supseteq H$ and $f\circ g(K)=K$. Hence, $Y_{1}$ and $\operatorname{X}(G/e,\tilde{l})$ have the same homotopy type (see Section 2.1). The graph $G/e$ has no $\tilde{l}$-separating edges (since $(G,l)$ has no $l$-separating edge), so by induction, we see that $Y_{1}\simeq\operatorname{X}(G/e,\tilde{l})$ is homotopy equivalent to a non-trivial wedge of $d(G/e,\tilde{l})$-spheres. Since $e$ is not a loop, we have $\operatorname{rk}(G/e)=\operatorname{rk}(G)$. We also have $k(\tilde{l})=k(l)$ because not both endpoints of $e$ are labelled. Hence, $d(G/e,\tilde{l})=d(G,l)$. ∎
Now we use this subposet to show the desired connectivity for $\operatorname{X}(G,l)$. If $Y_{1}=\operatorname{X}(G,l)$ then the result immediately follows from Claim 1. Thus we may assume that $G\setminus e\in\operatorname{X}(G,l)$. It follows that $\operatorname{X}(G,l)$ is obtained from $Y_{1}=\operatorname{X}(G,l)\setminus\left\{G\setminus e\right\}$ by attaching the star of $G\setminus e$ along $\operatorname{lk}(G\setminus e)$. That is, if $Y_{2}$ is the downward closure of $G\setminus e$ in $\operatorname{X}(G,l)$, then
$|\operatorname{X}(G,l)|=|Y_{1}|\cup_{|Y_{1}\cap Y_{2}|}|Y_{2}|,\text{ where }Y_{1}\cap Y_{2}=\operatorname{lk}(G\setminus e).$
The poset $Y_{2}$ is contractible as it has the unique maximal element $G\setminus e$, and $Y_{1}$ is homotopy equivalent to a wedge of $d(G,l)$-spheres by Claim 1. Hence, the following claim implies that $\operatorname{X}(G,l)$ is homotopy equivalent to a non-trivial wedge of $d(G,l)$-spheres and finishes the proof of Lemma 2.3.
@@
The poset $\operatorname{lk}(G\setminus e)$ is homotopy equivalent to a wedge of $(d(G,l)-1)$-spheres and either $\operatorname{lk}(G\setminus e)$ or $Y_{1}$ are non-contractible.
@@
The link $\operatorname{lk}(G\setminus e)$ consists of all $H\in\operatorname{X}(G,l)$ such that $H<G\setminus e$. It is isomorphic to $\operatorname{X}(G\setminus e,l^{\prime})$, where $l^{\prime}$ is the labelling induced by $l$. As $e$ is either non-separating or a leaf, $G\setminus e$ is connected. Hence, by induction, $\operatorname{X}(G\setminus e,l^{\prime})\cong\operatorname{lk}(G\setminus e)$ is homotopy equivalent to a wedge of $d(G\setminus e,l^{\prime})$-spheres and non-contractible if $G\setminus e$ contains no $l^{\prime}$-separating edge. By Claim 1, either this is the case or $Y_{1}$ is non-contractible. It remains to show that either $d(G\setminus e,l^{\prime})$ and $d(G,l)-1$ are equal or both are smaller than 00.
If $e$ is non-separating, then $\operatorname{rk}(G\setminus e)=\operatorname{rk}(G)-1$ and $k(l^{\prime})=k$. Hence the equality $d(G\setminus e,l^{\prime})=d(G,l)-1$ holds. Now assume that $e$ is a leaf. If $k\leq 1$ and $n=0$, then $\operatorname{X}(G\setminus e,l^{\prime})=\emptyset$. This is consistent with $d(G,l)=-2$. Otherwise, $(G,l)$ not having $l$-separating edges implies that the valence-one vertex of $e$ is labelled and that $k\geq 2$. Hence, we have $k(l^{\prime})=k-1\geq 1$ and $\operatorname{rk}(G\setminus e)=\operatorname{rk}(G)$, so we again get equality $d(G\setminus e,l^{\prime})=d(G,l)-1$. ∎
@@ 3. Proof of the main theorems
In this section, we prove Theorem 1.1, Theorem 1.3 and Theorem 1.2. We will frequently use Quillen’s fibre theorem as stated in [PW:Decompositions, Theorem 5.3] to show that certain maps are highly connected.
Throughout the section, we fix a finitely generated group $A$ and a (possibly empty) free factor system $\mathcal{A}$ of $A$. We set
@@
We denote by $\mathcal{FS}=\mathcal{FS}(A,\mathcal{A})$ the poset of free splittings of $A$ relative to $\mathcal{A}$. Its elements can be seen as pairs $S=(\mathbb{G},\mathfrak{m})$, where $\mathbb{G}$ is a graph of groups with underlying labelled graph a connected core graph in the sense of Section 2, and where $\mathfrak{m}\colon A\to\pi_{1}(\mathbb{G})$ is an isomorphism which is well-defined up to inner automorphisms. The poset relation is given by saying that $S>S^{\prime}$ if $S^{\prime}$ is obtained from $S$ by collapsing a subgraph. For $S\in\mathcal{FS}$, we denote by $\mathcal{V}(S)$ the free factor system given by the conjugacy classes of its vertex groups. By definition, we have $\mathcal{V}(S)\in\mathcal{FF}\cup\{\mathcal{A}\}$, i.e. $\mathcal{A}\sqsubseteq\mathcal{V}(S)$. We refer the reader to [Bru:buildingsfreefactora, Section 4.1] for more detailed definitions.
We set $L=L(A,\mathcal{A})=\{S\in\mathcal{FF}\mathrel{{\>:\>}}\mathcal{V}(S)=\mathcal{A}\}$ and $b\mathcal{FS}=\mathcal{FS}\setminus L$. Here, $L$ is the spine of relative Outer space $\mathcal{O}(A,\mathcal{A})$ in the sense of [GL:outerspacefree], and $b\mathcal{FS}$ is its simplicial boundary. This simplicial boundary has the homotopy type of the poset $\mathcal{FF}$ that we are interested in:
@@
This is part of [Bru:buildingsfreefactora, Corollary 4.35]. ∎
Now we prove that $p_{1}$ is highly connected. Section 4.6.4 of [Bru:buildingsfreefactora] studies this map for the case $A=F_{n}$ and $\mathcal{A}=\emptyset$. Here we use Lemma 2.3 to generalise this to an arbitrary pair $(A,\mathcal{A})$.
@@
Note that $e(n,k)$ agrees with $d(G,l)$ from Lemma 2.3 except if $k=1$. This case distinction will become clear in the following proof.
Proof of Lemma 3.3.
Let $S\in L$ and let $(G,l)$ be the underlying labelled graph. Note that $G$ has exactly $k$ labelled vertices and $\operatorname{rk}(G)=n$. By [Bru:buildingsfreefactora, Lemma 4.36], the fibre $p_{1}^{-1}(L_{\leq S})$ is homotopy equivalent to $p_{1}^{-1}(S)$. (The proof of [Bru:buildingsfreefactora, Lemma 4.36] is formulated for $A=F_{n}$, $\mathcal{A}=\emptyset$, but it generalises verbatim to the setup here.) The latter is exactly the poset $\operatorname{C}(G,l)$ defined in Section 2. This is $(d(G,l)-1)$-connected by Lemma 2.1 and Lemma 2.3. For $k\neq 1$, we have $d(G,l)=e(n,k)$ and the claim directly follows from Quillen’s fibre theorem [PW:Decompositions, Theorem 5.3].
For $k=1$, we have $d(G,l)=n-2$, so the same argument would give us one degree of connectivity less than claimed. Assume we are in this case. We will prove that
@@
[…]
$G$$v$$\widetilde{G}$$v$
@@
As a corollary, we get Theorem 1.2, which states that $\mathcal{FF}$ is $(n+k-4)$-connected for $k\geq 2$ and $(n-2)$-connected for $k\in\left\{0,1\right\}$:
Proof of Theorem 1.2.
As pointed out in the introduction, the case $k=0$, i.e. $\mathcal{A}=\emptyset$, only occurs if $A$ is a free group. In this case, $n=\operatorname{rk}(A)$ and Theorem 1.2 follows from [BG:Homotopytypecomplex, Theorem B]. Hence, we can assume that $k\neq 0$; we can also assume that $n+k\geq 2$ as there is nothing to show otherwise. Under these assumptions, the poset $L$ is contractible by [GL:outerspacefree, Section 4]. Thus by Lemma 3.1, Lemma 3.2 and Lemma 3.3, $\mathcal{FF}$ is $(n+k-4)$-connected for $k\geq 2$, and $(n-2)$-connected if $k=1$. In any case, $\mathcal{FF}$ is always at least $(n-2)$-connected. ∎
The following is a corollary of [HandelMosher, Lemma 2.5].
@@
We are now ready to prove Theorem 1.3, i.e. to show that $\mathcal{PB}$ is Cohen–Macaulay of dimension $n-1$.
Proof of Theorem 1.3.
@@
In view of the previous paragraph, it is enough to show that the inclusion map $i$ is an $(n-1)$-equivalence, and that the codomain is $(n-2)$-connected. To that end, we investigate the fibres and apply Quillen’s fibre theorem.
Let $\mathcal{A}^{\prime}\in\mathcal{FF}\setminus\mathcal{PB}$. By Proposition 3.4, we have
@@
Since a class $[B]\in\mathcal{A}^{\prime}$ gives a proper subgroup $B<A$, we have
@@
by Equation 1. Therefore, by induction, each $\mathcal{PB}(B,\mathcal{A}_{[B]})$ is $(\operatorname{crk}_{B}(\mathcal{A}_{[B]})-2)$-connected, and so the join of Equation 2 is
@@
On the other hand, since $\mathcal{A}\sqsubset\mathcal{A}^{\prime}$, by Theorem 1.2 we have that

### v2

@@ 2.3. Posets of core graphs
This is because every $H\in\operatorname{X}(G,l)$ contains a unique maximal core subgraph $\mathring{H}\in\operatorname{C}(G,l)$. The assignment $H\mapsto\mathring{H}$ defines a monotone poset map $\operatorname{X}(G,l)\to\operatorname{C}(G,l)$ such that $H\geq\mathring{H}$ and restricts to the identity on $\operatorname{C}(G,l)$. Hence $\operatorname{X}(G,l)$ deformation retracts to $\operatorname{C}(G,l)$ (see [4, Corollary 2.5]). ∎
In [4, Proposition 4.14], it was shown that for $k(l)=0$, the poset $\operatorname{X}(G,l)$ is contractible if and only if $G$ has a separating edge. The following definition is set up in such a way that this result generalises to the cases $k(l)>0$.
@@
Remark 2.3.
For $k\leq 1$, an edge is $l$-separating if and only if it is separating and $G$ has non-trivial fundamental group. In particular, the $l$-separating edges of $(G,l)$ are the same as the $\emptyset$-separating edges of the trivially labelled graph $(G,\emptyset)$.
@@
To explain the case distinction, note that if $k(l)=1$, then $\operatorname{X}(G,l)=\operatorname{X}(G,\emptyset)$. In [4, Proposition 4.14], Lemma˜2.4 was deduced from [24, Proposition 2.2] for the case $k=0$, $n\geq 2$ under the assumption that every vertex of $G$ has valence at least three.
Proof of Lemma˜2.4
Suppose first that $G$ contains an $l$-separating edge $e$. We want to show that $\operatorname{X}(G,l)$ is contractible. Let $H\in\operatorname{X}(G,l)$, i.e. it has at least one connected component that has trivial fundamental group or contains at least two labelled vertices. We claim that the same is true for $H\setminus\left\{e\right\}$. If $e\not\in E(H)$, this is clear. So assume $e\in E(H)$. Note that $e\not\in\operatorname{X}(G,l)$, so $H\setminus\left\{e\right\}\neq\emptyset$. If $H$ has a connected component with non-trivial fundamental group, the same is true for $H\setminus\left\{e\right\}$ as $e$ is separating in $H$. On the other hand, if $H$ has a component with at least two labelled vertices, then by Item˜2 of Definition˜2.2, the same is true for $H\setminus\left\{e\right\}$. Hence $H\setminus\left\{e\right\}$ is an element of $\operatorname{X}(G,l)$, as claimed. Using this, the assignment $H\mapsto H\setminus\left\{e\right\}$ gives a monotone poset map $\operatorname{X}(G,l)\to\operatorname{X}(G,l)$, so a homotopy equivalence to its image. Using Item˜3 of Definition˜2.2, $G\setminus e$ is an element of $\operatorname{X}(G,l)$, so it lies in this image. It clearly forms a maximal element, so the image and hence $\operatorname{X}(G,l)$ are contractible.
Now assume that $G$ contains no $l$-separating edges. We prove that $\operatorname{X}(G,l)$ is a non-trivial wedge of $d(G,l)$-spheres by induction on the number of edges of $G$. If $G$ has just a single edge, then this edge is not $l$-separating and $\operatorname{X}(G,l)$ is empty, so in particular non-contractible (this is true independently of whether the edge forms a loop and whether it has labelled endpoints or not). Hence, the claim is true. In what follows, we will assume that there are at least two edges.
@@
Suppose that $G$ contains a loop $e$. We claim that $G\setminus e$ is an element of $\operatorname{X}(G,l)$. As $G$ has at least two edges, $G\setminus e$ is non-empty. If $G\setminus e$ was a tree with only one labelled vertex, then also $G$ has just one labelled vertex and all edges except $e$ are separating. But by Remark˜2.3, they are $l$-separating, which contradicts our assumption. Hence, $G\setminus e$ either has non-trivial fundamental group or at least two labelled vertices, which proves that $G\setminus e\in\operatorname{X}(G,l)$.
Using this, the poset $\operatorname{X}(G,l)$ can be written as the union of two subposets $X_{1}$ and $X_{2}$ as follows. Let
@@
where $l^{\prime}$ is the labelling induced by $l$. This subposet has $G\setminus e$ as its unique maximal element and hence is contractible. The second subposet is given by $\operatorname{X}(G\setminus e,l^{\prime})$ together with all proper subgraphs of $G$ containing $e$ (such a subgraph is always contained in $\operatorname{X}(G,l)$ because it contains the loop $e$ and hence has non-trivial fundamental group). That is,
@@
By the previous paragraph, we can now assume that $G$ contains no loops. If $G$ is a tree, let $e$ be a leaf of $G$. Otherwise, let $e$ be a non-separating (and non-loop) edge. Note that in either case, $G\setminus e$ is connected. We will prove the two cases in parallel. Let $Y_{1}:=\operatorname{X}(G,l)\setminus\left\{G\setminus e\right\}$ and denote by $(G\setminus e,l^{\prime})$ the labelled graph obtained from $G$ by removing $e$.
@@
To prove this, we consider two cases. The first is that $e$ has both endpoints labelled. We claim that in this case, $Y_{1}$ is contractible and different from $\operatorname{X}(G,l)$. As $G$ has more than one edge and both endpoints of $e$ are labelled, we have $\left\{e\right\}\in Y_{1}$. We get a monotone map $Y_{1}\to Y_{1}$ by the assignment $H\mapsto H\cup\left\{e\right\}$. Its image has a unique minimal element $\left\{e\right\}$, so $Y_{1}$ is contractible. It remains to show that $Y_{1}$ is different from $\operatorname{X}(G,l)$, i.e. that $G\setminus e\in\operatorname{X}(G,l)$. Recall that $G\setminus e$ is connected and that either $G$ is a tree with $e$ a leaf or $e$ is non-separating. If $e$ is non-separating, then $G\setminus e$ still contains both (labelled) endpoints of $e$. Hence, we have indeed $G\setminus e\in\operatorname{X}(G,l)$. If, on the other hand, $G$ is a tree and $e$ is a leaf, the assumption that $G$ has no $l$-separating edge implies that there must be at least one labelled vertex that is not an endpoint of $e$. Hence, $G\setminus e$ is a connected graph with at least two labelled vertices and we again have $G\setminus e\in\operatorname{X}(G,l)$.
[…]
that sends a subgraph to its image in the quotient $G/e$. We define a poset map in the other direction as follows. Let $v_{e}$ be the vertex of $G/e$ to which $e$ was collapsed. Then define
[…]
$K$ $\mapsto\begin{cases}K\cup\left\{e\right\}&v_{e}\in V(K),\\ K&\text{otherwise}.\end{cases}$
[…]
Now we use this subposet to show the desired connectivity for $\operatorname{X}(G,l)$. If $Y_{1}=\operatorname{X}(G,l)$ then the result immediately follows from Claim 1. Thus we may assume that $G\setminus e\in\operatorname{X}(G,l)$. It follows that $\operatorname{X}(G,l)$ is obtained from $Y_{1}=\operatorname{X}(G,l)\setminus\left\{G\setminus e\right\}$ by attaching the star of $G\setminus e$ along its link $\operatorname{lk}(G\setminus e)$. That is, if $Y_{2}$ is the downward closure of $G\setminus e$ in $\operatorname{X}(G,l)$ (i.e. the lower interval $\operatorname{X}(G,l)_{\subseteq G\setminus e}$), then
$|\operatorname{X}(G,l)|=|Y_{1}|\cup_{|Y_{1}\cap Y_{2}|}|Y_{2}|,\text{ where }Y_{1}\cap Y_{2}=\operatorname{X}(G,l)_{\subsetneq G\setminus e}.$
Here, $\operatorname{lk}(G\setminus e)$ is the order complex of $\operatorname{X}(G,l)_{\subsetneq G\setminus e}$. The poset $Y_{2}$ is contractible as it has the unique maximal element $G\setminus e$, and $Y_{1}$ is homotopy equivalent to a wedge of $d(G,l)$-spheres by Claim 1. Hence, the following claim implies that $\operatorname{X}(G,l)$ is homotopy equivalent to a non-trivial wedge of $d(G,l)$-spheres and finishes the proof of Lemma˜2.4.
@@
The poset $\operatorname{X}(G,l)_{\subsetneq G\setminus e}$ is homotopy equivalent to a wedge of (d(G,l)1)(d(G,l)-1)-spheres and either $\operatorname{X}(G,l)_{\subsetneq G\setminus e}$ or $Y_{1}$ are non-contractible.
@@
Note that $\operatorname{X}(G,l)_{\subsetneq G\setminus e}$ is isomorphic to $\operatorname{X}(G\setminus e,l^{\prime})$, where $l^{\prime}$ is the labelling induced by $l$. Since $G\setminus e$ is connected, by induction, $\operatorname{X}(G\setminus e,l^{\prime})$ is homotopy equivalent to a wedge of $d(G\setminus e,l^{\prime})$-spheres, and it is not contractible if $G\setminus e$ contains no $l^{\prime}$-separating edge. By Claim 1, either this is the case or $Y_{1}$ is non-contractible. It remains to show that either $d(G\setminus e,l^{\prime})$ and d(G,l)1d(G,l)-1 are equal or both are smaller than 0.
Here, we need to distinguish the cases where $e$ is either non-separating or a leaf. If $e$ is non-separating, then $\operatorname{rk}(G\setminus e)=\operatorname{rk}(G)-1$ and $k(l^{\prime})=k$. Hence the equality $d(G\setminus e,l^{\prime})=d(G,l)-1$ holds. Now assume that $e$ is a leaf. If $k\leq 1$ and $n=0$, then $\operatorname{X}(G\setminus e,l^{\prime})=\emptyset$. This is consistent with d(G,l)=2d(G,l)=-2. Otherwise, $(G,l)$ not having $l$-separating edges implies that the valence-one vertex of $e$ is labelled and that $k\geq 2$. Hence, we have $k(l^{\prime})=k-1\geq 1$ and $\operatorname{rk}(G\setminus e)=\operatorname{rk}(G)$, so we again get equality $d(G\setminus e,l^{\prime})=d(G,l)-1$. ∎
@@ 3. Proof of the main theorems
In this section, we prove Theorem˜1.1, Theorem˜1.3 and Theorem˜1.2. We will frequently use Quillen’s fibre theorem as stated in [20, Theorem 5.3] to show that certain maps are highly connected (that is, they induce isomorphism between the homotopy groups up to a large enough degree).
Throughout the section, we fix a finitely generated group $A$ and a (possibly empty) free factor system $\mathcal{A}$ of $A$ , and write $\mathcal{FF}=\mathcal{FF}(A,\mathcal{A})$. We set
@@
We denote by $\mathcal{FS}=\mathcal{FS}(A,\mathcal{A})$ the poset of free splittings of $A$ relative to $\mathcal{A}$. Its elements can be seen as pairs $S=(\mathbb{G},\mathfrak{m})$, where $\mathbb{G}$ is a graph of groups with underlying labelled graph a connected core graph in the sense of Section˜2, and where $\mathfrak{m}\colon A\to\pi_{1}(\mathbb{G})$ is an isomorphism which is well-defined up to inner automorphisms. The poset relation is given by saying that $S>S^{\prime}$ if $S^{\prime}$ is obtained from $S$ by collapsing a subgraph. For $S\in\mathcal{FS}$, we denote by $\mathcal{V}(S)$ the free factor system given by the conjugacy classes of its vertex groups. By definition, we have $\mathcal{V}(S)\in\mathcal{FF}\cup\{\mathcal{A}\}$, i.e. $\mathcal{A}\sqsubseteq\mathcal{V}(S)$. We refer the reader to [4, Section 4.1] for more detailed definitions.
We set $L=L(A,\mathcal{A})=\{S\in\mathcal{FF}\mathrel{{\>:\>}}\mathcal{V}(S)=\mathcal{A}\}$ and $b\mathcal{FS}=\mathcal{FS}\setminus L$. Here, $L$ is the spine of relative Outer space $\mathcal{O}(A,\mathcal{A})$ in the sense of [12], and $b\mathcal{FS}$ is its simplicial boundary. This simplicial boundary has the homotopy type of the poset $\mathcal{FF}$ that we are interested in:
@@
This is part of [4, Corollary 4.35]. ∎
Now we prove that $p_{1}$ is highly connected. Section 4.6.4 of [4] studies this map for the case $A=F_{n}$ and $\mathcal{A}=\emptyset$. Here we use Lemma˜2.4 to generalise this to an arbitrary pair $(A,\mathcal{A})$.
@@
Note that $e(n,k)$ agrees with $d(G,l)$ from Lemma˜2.4 except if $k=1$. This case distinction will become clear in the following proof.
Proof of Lemma˜3.3.
Let $S\in L$ and let $(G,l)$ be the underlying labelled graph. Note that $G$ has exactly $k$ labelled vertices and $\operatorname{rk}(G)=n$. By [4, Lemma 4.36], the fibre $p_{1}^{-1}(L_{\leq S})$ is homotopy equivalent to $p_{1}^{-1}(S)$. (The proof of [4, Lemma 4.36] is formulated for $F_{n}, 𝒜=\mathcal{A}=\emptyset$, but it generalises verbatim to the setup here.) The latter is exactly the poset $\operatorname{C}(G,l)$ defined in Section˜2. This is (d(G,l)1)(d(G,l)-1)-connected by Lemma˜2.1 and Lemma˜2.4. For $k\neq 1$, we have $d(G,l)=e(n,k)$ and the claim directly follows from Quillen’s fibre theorem [20, Theorem 5.3].
For $k=1$, we have d(G,l)=n2d(G,l)=n-2, so the same argument would give us one degree of connectivity less than claimed. Assume we are in this case. We will prove that
@@
is (n2)(n-2)-connected for all $S\in L$. Applying Quillen’s fibre theorem then gives us the slightly stronger connectivity statement. Note that $n\neq 0$ since $\mathcal{A}\neq\{[A]\}$. If $L_{>S}$ is non-empty (i.e. $S$ is not maximal), then $p_{1}^{-1}(L_{\leq S})*L_{>S}$ is at least (n2)(n-2)-connected. Suppose now that $S$ is maximal. We prove that this implies that $G$ contains an $l$-separating edge. Since $G$ has only one labelled vertex, say $v$, and $\operatorname{rk}(G)=n\neq 0$, it is enough to show that $v$ has valence one (and hence the $l$-separating edge is the leaf containing $v$). Indeed, if $v$ has valence at least two, then we can “push away” the labelling from $v$ by adding a new edge with one endpoint $v$ and the other endpoint containing the labelling that $v$ had in $G$ (and thus $v$ is now unlabelled). See Figure˜2. This graph defines an element of $L$ that lies in $L_{>S}$ (all unlabelled vertices are at least trivalent), contradicting the maximality of $S$. Therefore $v$ must be a vertex of valence one. Thus $G$ contains an $l$-separating edge, and by Lemma˜2.4, $p_{1}^{-1}(L_{\leq S})\simeq\operatorname{C}(G,l)\simeq\operatorname{X}(G,l)$ is contractible. In particular this implies that $p_{1}^{-1}(L_{\leq S})*L_{>S}$ is also (n2)(n-2)-connected. This concludes the proof. ∎
$G~\widetilde{G}vv$
@@
As a corollary, we get Theorem˜1.2, which states that $\mathcal{FF}$ is (n+k4)(n+k-4)-connected for $k\geq 2$ and (n2)(n-2)-connected for $k\in\left\{0,1\right\}$:
Proof of Theorem˜1.2.
As pointed out in the introduction, the case $k=0$, i.e. $\mathcal{A}=\emptyset$, only occurs if $A$ is a free group. In this case, $n=\operatorname{rk}(A)$ and Theorem˜1.2 follows from [6, Theorem B]. Hence, we can assume that $k\neq 0$; we can also assume that $n+k\geq 2$ as there is nothing to show otherwise. Under these assumptions, the poset $L$ is contractible by [12, Section 4]. Thus by Lemma˜3.1, Lemma˜3.2 and Lemma˜3.3, $\mathcal{FF}$ is (n+k4)(n+k-4)-connected for $k\geq 2$, and (n2)(n-2)-connected if $k=1$. In any case, $\mathcal{FF}$ is always at least (n2)(n-2)-connected. ∎
The following is a corollary of [13, Lemma 2.5].
@@
We are now ready to prove Theorem˜1.3, i.e. to show that $\mathcal{PB}:=\mathcal{PB}(A,\mathcal{A})$ is Cohen–Macaulay of dimension n1n-1.
Proof of Theorem˜1.3.
@@
In view of the previous paragraph, it is enough to show that the inclusion map $i$ is an (n1)(n-1)-equivalence, and that the codomain is (n2)(n-2)-connected. To that end, we investigate the fibres and apply Quillen’s fibre theorem.
Let $\mathcal{A}^{\prime}\in\mathcal{FF}\setminus\mathcal{PB}$. By Proposition˜3.4, we have
@@
Since a class $B]\in\mathcal{A}^{\prime}$ gives a proper subgroup $B<A$, we have
@@
by Equation˜1. Therefore, by induction, each $\mathcal{PB}(B,\mathcal{A}_{[B]})$ is $(\operatorname{crk}_{B}(\mathcal{A}_{[B]})-2)$-connected, and so the join of Equation˜2 is
@@
On the other hand, since $\mathcal{A}\sqsubset\mathcal{A}^{\prime}$, by Theorem˜1.2 we have that